How many ways can 6 students sit in 3 seats in a row ?
Found 2 solutions by ewatrrr, sudhanshu_kmr:
Answer by ewatrrr(24785)

Hi,
The number of combinations of a set of n objects taken r at a time is given by nCr = (n!)/(r!(n - r)!).
.
6 Choose 3!
.
6C3 =(6!)/(3!(6 - 3)! = 6*5*4/ 3*2 = 20 ways

Answer by sudhanshu_kmr(1152)


No.of ways to select 3 student from 6 student to sit in 3 seats = 6C3 = 20

no. of ways to arrange these 3 student on seat = 3!
= 6

total no. of ways to = 20 *6 = 120